OpenStudy (anonymous):

This season, Lisa's lacrosse team has won 2/3 of their home games (games played at Lisa's school), but just 2/5 of their away games (games played at other schools). In total, Lisa's team has won 26 games out of 49 games they have played. How many home games has Lisa's team played? Explain how you solved the problem. MEDAL ANYONE WHO SOLVES

OpenStudy (anonymous):

Good luck little brother :)

OpenStudy (kropot72):

Let the number of away games played be a, and let the number of home games played be h. Now we can form two equations from the given information: \[a+h=49\ ...............(1)\] \[\frac{2}{5}a+\frac{2}{3}h=26\ ...........(2)\] Just solve the system of equations (1) and (2) to find h.

OpenStudy (kropot72):

Multiplying both sides of (2) by 5/2 gives \[a+\frac{10}{6}h=65\ ...........(3)\] and subtracting equation(1) from equation (3) results in \[\frac{4}{6}h=16\ ..........(4)\] Can you solve equation (4) to find h?